A motorist travels for 3okm at xkm/h and for 9okm at ykm/h and takes place 2.5hours for the journey. if the speed are interchanging, the journey take 2whole number 1/6 find x and y
Answer by mananth(16949)

Distance = speed * time
Time = distance/speed
t=d/r
First condition
Time Part I + time part II = 2.5 hour
30/x+90/y = 2.5
Let 1/x=a & 1/y = b
30a+90b=5/2.......................(1)
Second condition
90/x+30/y = 2 1/6 = 13/6
90a+30b= 13/6......................(2)
add (1) & (2)
120a+120b=5/2+13/6
120a+120b=14/3
divide by 120
a+b= 14/360.............(3)
(1)-(2)
-60a+60b=1/3
divide by -60
a-b=-1/180...................(4)
(3)+(4)
2a= 14/360-1/180
a=7/360 -1/360
a= 6/360
a=1/60
1/x= 1/60
x= 60

a+b = 14/360
b= 14/360 -1/60
b=8/360
b=1/y = 8/360
y= 360/8
y=45